# Variability of MHD Instabilities in Benign Termination of High-Current Runaway Electron Beams in the JET and DIII-D Tokamaks

### v1

@@ 3.3 Phenomenology of the terminating MHD events on JET and DIII-D
Figure 8: Internal inductance $l_{i}$ versus the edge safety factor $q_{\text{edge}}$ immediately before the terminating MHD event on JET (Panel a) and DIII-D (Panel b). Error bars reflect the systematic uncertainty on the $l_{i}$ measurement, mainly depending on the elongation $\kappa$, as discussed in the corresponding Section 3.3 and Appendix A.1. The black dotted lines recreate the empirical MHD stability boundary reported by Wesson et al. [Wesson_1989]. The encircled data point in Panel (a) corresponds to discharge #95135, which was previously used as a reference case for benign termination at JET and is found to be nearly singular.
@@
combined $\tau_{A}\cdot\gamma$ $\approx 0.003$ $\approx 0.0015$
@@ 4 MHD modeling of the termination boundaries
As the internal inductance and edge safety factor were found in the experiment to determine the MHD stability boundaries, both were varied in the following scans to track their effect. To generate otherwise self-similar equilibria, the NEMEC equilibrium solver was used [HIRSHMAN_NEMEC]. As an example, the resulting current and safety-factor profiles are shown in Fig. 10(a) and (b) as a function of the square root of the normalized toroidal flux $\rho_{\text{tor}}$. The depicted $l_{i}$ values are representative of cases just beyond the upper and lower stability boundaries found for $q_{\text{edge}}=3.2$, with the flat current profile for low $l_{i}=0.65$ (dashed blue line) and the more peaked one for $l_{i}=1.49$ (solid red line). The corresponding instability eigenfunctions from the MHD modeling in Panel (c) suggest strong, external kink modes (peaking outside mid-radius) for low $l_{i}$ values and weaker, internal kink modes (peaking inside mid-radius) for higher $l_{i}$ values.
@@
Finding internal kink modes seems to be a trivial result, given that the core safety factors drop below $1$, see Panel (b). Therefore, to improve the understanding of the upper stability boundary, another scan of current profiles was performed in which the safety factor was enforced to remain larger than $1$ for high $l_{i}$ values. In this scan, the internal kink disappeared and only a small, underlying, external resistive mode was found, without a clear instability boundary. Therefore, in the linear, resistive picture, the existence of an upper boundary, as observed in the experiments, relies necessarily on the assumption that the core safety factors can fall below $1$. Internal kink modes are assumed to serve as a seed for (resistive) non-linear events, resulting in stochastization and, hence, deconfinement of the RE beam. Modeling such mode interactions would require nonlinear MHD calculations, which are the subject of future work and cannot be applied for scanning large parameter spaces due to their computational expense.
@@
Figure 10: Examples from the CASTOR3D modeling, varying the current density (Panel a) for fixed edge values of the safety-factor profile (Panel b) to scan different $l_{i}$ values. The $l_{i}$ values for which the curves are shown are just beyond the upper, respectively lower stability boundary. As can be seen in Panel (c), the mode eigenfunctions suggest different instability natures, with strong, external kink modes for low $l_{i}$ values (blue dashed line) and weaker, internal kink modes for high $l_{i}$ values (red solid line). A resistivity of $10^{-5}\penalty 10000\ \Omega\,$m is used.
@@
In addition, the effect of using different resistivities in the MHD calculations was investigated. For this purpose, the linear, normalized growth rates $\gamma\,\tau_{A}$ were scanned for the DIII-D reference case over a range of $l_{i}$ values, as shown in Fig. 11. One can see that using different resistivities modifies the calculated growth rates only for the high $l_{i}$ values identified with internal kink modes (red region). However, no new types of instabilities nor shifts of the stability boundaries are observed. In this linear picture, the external–kink events (blue region) are significantly more rapid in nature, yielding larger growth rates. A rather stable region is found for medium $l_{i}$ values (see yellow region), where normalized growth rates are below 1%, and in which stabilizing effects, such as flows or finite Larmor radius effects, can suppress such slowly growing modes. While this analysis is based on the growth rates, a realistic calculation of $\delta B$ values would require more sophisticated, non-linear MHD modeling with complex synthetic diagnostics and is the subject of future work. The vertical lines in Fig. 11 correspond to the $l_{i}$ values used in Fig. 10 and shown in the corresponding linestyle and color.

### v2

@@ 3.3 Phenomenology of the terminating MHD events on JET and DIII-D
Figure 9: Internal inductance $l_{i}$ versus the edge safety factor $q_{\text{edge}}$ immediately before the terminating MHD event on JET (Panel a) and DIII-D (Panel b). Error bars reflect the systematic uncertainty on the $l_{i}$ measurement, mainly depending on the elongation $\kappa$, as discussed in the corresponding Section 3.3 and Appendix A.1. The black dotted lines recreate the empirical MHD stability boundary reported by Wesson et al. [Wesson_1989]. The encircled data point in Panel (a) corresponds to discharge #95135, which was previously used as a reference case for benign termination at JET and is not found to be representative of the broader dataset.
@@
combined $\tau_{A}\cdot\gamma$ $\approx 0.003$ $\approx 0.015$
@@ 4 MHD modeling of the termination boundaries
As the internal inductance and edge safety factor were found in the experiment to determine the MHD stability boundaries, both were varied in the following scans to track their effect. To generate otherwise self-similar equilibria, the NEMEC equilibrium solver was used [HIRSHMAN_NEMEC]. As an example, the resulting current and safety-factor profiles are shown in Fig. 11(a) and (b) as a function of the square root of the normalized toroidal flux $\rho_{\text{tor}}$. The depicted $l_{i}$ values are representative of cases just beyond the upper and lower stability boundaries found for $q_{\text{edge}}=3.2$, with the flat current profile for low $l_{i}=0.65$ (dashed blue line) and the more peaked one for $l_{i}=1.49$ (solid red line). The corresponding instability eigenfunctions from the MHD modeling in Panel (c) suggest strong, external kink modes (peaking outside mid-radius) for low $l_{i}$ values and weaker, internal kink modes (peaking inside mid-radius) for higher $l_{i}$ values. The internal kink is numerically identified as an $m/n=1/1$ mode, while the external kink is identified as an $m/n=4/1$ mode.
@@
Finding internal kink modes seems to be a trivial result, given that the core safety factors drop below $1$, see Fig. 11(b). Therefore, to improve the understanding of the upper stability boundary, another scan of current profiles was performed in which the safety factor was enforced to remain larger than $1$ for high $l_{i}$ values. In this scan, the internal kink disappeared and only a small, underlying, external resistive mode was found, without a clear instability boundary. Therefore, in the linear, resistive picture, the existence of an upper boundary, as observed in the experiments, relies necessarily on the assumption that the core safety factors can fall below $1$. Owing to the absence of neutral beam injection and the Motional Stark Effect (MSE) diagnostic, this modeling result cannot currently be validated experimentally. However, it represents the most plausible scenario, as the experimentally observed upper kink stability cannot otherwise be reproduced by the MHD modeling. Internal kink modes are assumed to serve as a seed for (resistive) non-linear events, resulting in stochastization and, hence, deconfinement of the RE beam. Modeling such mode interactions would require nonlinear MHD calculations, which are the subject of future work and cannot be applied for scanning large parameter spaces due to their computational expense.
@@
Figure 11: Examples from the CASTOR3D modeling, varying the current density (Panel a) for fixed edge values of the safety-factor profile (Panel b) to scan different $l_{i}$ values. The $l_{i}$ values for which the curves are shown are just beyond the upper, respectively lower stability boundary. As can be seen in Panel (c), the mode eigenfunctions suggest different instability natures, with strong, external kink modes ($m/n=4/1$) for low $l_{i}$ values (blue dashed line) and weaker, internal kink modes ($m/n=1/1$) for high $l_{i}$ values (red solid line). A resistivity of $10^{-5}\penalty 10000\ \Omega\,$m is used.
@@
In addition, the effect of using different resistivities in the MHD calculations was investigated. For this purpose, the linear, normalized growth rates $\gamma\,\tau_{A}$ were scanned for the DIII-D reference case over a range of $l_{i}$ values, as shown in Fig. 12. One can see that using different resistivities modifies the calculated growth rates only for the high $l_{i}$ values identified with internal kink modes (red region). However, no new types of instabilities nor shifts of the stability boundaries are observed. In this linear picture, the external–kink events (blue region) are significantly more rapid in nature, yielding larger growth rates. A rather stable region is found for medium $l_{i}$ values (see yellow region), where normalized growth rates are below 1%, and in which stabilizing effects, such as flows or finite Larmor radius effects, can suppress such slowly growing modes. While this analysis is based on the growth rates, a realistic calculation of $\delta B$ values would require more sophisticated, non-linear MHD modeling with complex synthetic diagnostics and is the subject of future work. In addition, such future work could explore whether flatter current profiles and, consequently, flatter $q$-profiles could lead to more global mode structure, beneficial for efficient stochastication. A direct comparison with the growth rates assessed experimentally, e.g., in Tab. 1 or Fig. 10, shows that the predicted mode growth is faster than what is observed experimentally. This suggests that there are effects damping the mode growth in reality, such as nonlinear resistive effects or stabilizing influences from the wall, which are not included in the current modeling. The vertical lines in Fig. 12 correspond to the $l_{i}$ values used in Fig. 11 and shown in the corresponding linestyle and color.